Maggie's brother is 3 years younger than twice her age. The sum of their ages is 24. How old is Maggie.
Answer by ikleyn(52781)

.
Let x be the Maggie age.


Then the brother's age is (2x-3) years.


The condition says


x + (2x-3) = 24,    or


x + 2x - 3 = 24,


3x = 24 + 3,


3x = 27,


x = 27/3 = 9.


Answer.  Maggie is 9 years old.

Solved.